# Conceptualization, Methodology, Writing – review & editing, Supervision

### v3

@@ Abstract
Pedestrian safety is an increasing concern, particularly near bus stops along urban transit corridors where pedestrian-vehicle interactions frequently occur. Previous research has predominantly analyzed pedestrian crashes at intersections and midblock locations, treating bus stops mainly as contributing factors rather than distinct safety assessment sites. Consequently, there is a notable gap in understanding the specific influence of factors such as traffic exposure, roadway environment, and bus stop design on pedestrian crash frequency near bus stops. Additionally, conventional statistical models often neglect unobserved site-specific variations, potentially leading to inaccurate identification of high-risk locations and ineffective safety interventions. To address these critical gaps, this study develops a comprehensive framework explicitly focusing on pedestrian safety near bus stops. The proposed framework utilizes the Random Parameters Negative Binomial-Lindley (RPNB-L) statistical model, effectively capturing unobserved heterogeneity and site-specific variations. Using data from 596 bus stops in Fort Worth, Texas (2018–2022), results show that higher pedestrian crash frequencies are significantly associated with increased AADT, greater pedestrian boarding activities, and the absence of critical safety features such as crosswalks, medians, and adequate lighting. Further analysis indicates increased crash risks near schools and on roads with lower speed limits ($\leq$35 mph) while signalized intersections, far-side bus stops, mixed-use developments, and sidewalks are associated with reduced crash risks. To enhance safety assessment, the study also applies a full Bayes (FB) based potential for safety improvement (PSI) metric, systematically identifying hazardous stops and high-risk corridors to help prioritize safety interventions. By integrating statistical count modeling with proactive risk assessment, this study bridges critical research gaps and provides data-driven insights to improve pedestrian safety near bus stops.
@@ 1 Introduction
Accordingly, the main objective of this study is to investigate the contributing factors to pedestrian crashes near bus stops in the United States that account for traffic exposure, roadway characteristics, and bus stop design. More specifically, this research (i) develops a comprehensive pedestrian crash analysis framework near bus stops along urban transit corridors, incorporating factors such as traffic exposure, roadway environment, and bus stop design, (ii) applies a random parameters approach to capture unobserved heterogeneity by site-specific variations, and (iii) employs a full bayes (FB)–based Potential for Safety Improvement (PSI) metric to identify and prioritize high-risk transit corridors. By addressing notable gaps in the existing literature, this work offers a systematic approach to understanding and mitigating pedestrian crashes near bus stops.
To accomplish these goals, the study integrates five years (2018–2022) of crash records from 596 bus stop locations in Fort Worth, Texas, with complementary data on roadway conditions, traffic exposure, and bus stop design. The traffic exposure data was captured from the agency’s database, and the roadway environment and bus stop design information was collected from the Rhino database and Google Earth & Street View. To account for site-specific variations and unobserved heterogeneity in crash frequency among different bus stops, this study adopts the RPNB-L as the primary modeling approach, while baseline models (NB-L, and RPNB-GE) are also estimated for comparative assessment. Additionally, the FB-based PSI metric is applied to systematically identify hazardous bus stops and transit corridors, enabling a proactive approach to pedestrian safety interventions. By bridging these critical research gaps, this study introduces a novel framework for evaluating and mitigating pedestrian crash risks near bus stops, contributing to the advancement of urban transit safety.
@@ 3.1 Negative Binomial-Lindley (NB-L)
[…]
@@
Where $\mu_{i}$=$\exp(b_{0}+\sum_{j=1}^{q}b_{j}X_{j})$ represents the expected number of crashes, and $E(\lambda_{i})$ = $\frac{\theta+2}{\theta(\theta+1)}$ is the expected value of the Lindley-distributed term. Replacing the values for $\mu_{i}$ and $E(\lambda_{i})$, the mean response function becomes:
@@ 4 Data description
This section provides an overview of this study’s data sources and variables. Subsection 4.1 introduces the pedestrian crash data obtained from the Texas Department of Transportation’s (TxDOT) Crash Records Information System (CRIS) database. Subsection 4.2 presents the additional information collected on bus stop design, missing roadway attributes, and exposure-related features. This complementary data enhances the historical crash analysis, offering a more comprehensive understanding of factors affecting pedestrian safety near bus stops.
@@ 4.1 Pedestrian crash data
The study utilized bus stop data from the Fort Worth-area transit service, Trinity Metro, which was accessed through their official website https://www.trinitymetro.org. Trinity Metro provided Geographic Information Systems (GIS) shapefiles containing detailed information on bus routes and stops, including amenities and wheelchair accessibility. The dataset comprised approximately 2,500 bus stops, each identified by the main street and a nearby cross-street. Additionally, Trinity Metro provided two spreadsheets containing boarding and alighting data categorized by stop location and day type (weekday, Saturday, and Sunday) for two distinct periods: pre-pandemic (September 2019) and during the pandemic (September 2021).
Pedestrian crash data for the years 2018 to 2022 were obtained from the Texas Department of Transportation’s (TxDOT) Crash Records Information System (CRIS). The dataset included only reportable crashes, which involved fatalities, injuries, or property damage exceeding \$1,000 and occurred on public roadways. To associate pedestrian crashes with bus stop locations, a spatial buffer of 250 feet was created around each bus stop. Only crashes that occurred within this radius and directly on the roadway associated with the bus stop were included. Crashes that fell within this radius but occurred on adjacent cross-streets were manually reviewed and excluded from the analysis to maintain data accuracy. Given the practical limitations associated with collecting detailed geometric and operational data for every bus stop, this study adopted a strategic sampling approach to facilitate a comprehensive analysis. Specifically, bus stops were selected based on two clearly defined criteria to ensure the representation all locations:
@@
    All bus stops with recorded pedestrian crashes.
@@
    Bus stops with reasonable pedestrian activity but no recorded crashes.
The first criterion captures locations with documented pedestrian crashes during the study period. Analyzing these sites enables a reactive investigation of associated factors with crash occurrences. In contrast, the second criterion focuses on bus stops that experienced reasonable pedestrian activity but no reported crashes. These sites were intentionally selected to allow for a proactive assessment of possible risk factors, as high pedestrian exposure - despite the absence of crashes—may still indicate latent safety concerns. By including both types of bus stops, the study aims to contrast hazardous and nonhazardous locations, thereby identifying overrepresented factors at crash-prone sites and potential protective features at safer ones. This dual approach strengthens the analysis by providing a comprehensive understanding of pedestrian safety dynamics near bus stops, grounded in both historical crash data and exposure-based risk anticipation.
Fig. 4.1 provides a detailed flowchart summarizing the data preparation process, including the integration of crash data with bus stop and roadway information. Based on the selection criteria described, 596 bus stops were ultimately chosen for detailed analysis, as illustrated in Fig. 4.1. Further, the study integrated pedestrian crash data with additional roadway geometric and traffic characteristics obtained from TxDOT’s 2019 Roadway Highway Inventory Network Offload (RHiNo) database. This comprehensive database includes location-specific roadway information of traffic and geometric data. This database also contains details on roadway design and traffic characteristics for both state and local roads, including historical traffic volumes from the past ten years.
@@
The RHiNo database, provided in GIS format, enabled seamless integration with bus stop information. It includes essential data, such as roadway segment location, system type, facility type, design characteristics (e.g., number of lanes, surface width, shoulder width, and median type), and current and historical traffic volumes. When specific geometric information (e.g., sidewalk presence, crosswalk availability, lighting conditions) was missing from the RHiNo database, the study supplemented this data using Google Earth® and Google Street View®. These tools proved to be valuable resources for collecting additional geometric information.
@@
In total, the pedestrian crash dataset included 448 pedestrian crashes, categorized according to severity levels. Among these incidents, 46 crashes (10.3%) involved fatalities (K), 394 crashes (87.8%) involved varying degrees of injury (levels A, B, or C), and the remaining 8 crashes (less than 2%) were classified as property damage only (PDO). Fig. 4.1 shows the yearly distribution of pedestrian crashes by severity level, while Fig.4.1 illustrates their general distribution characteristics. The data underscore significant pedestrian safety risks, highlighting the critical need for targeted safety enhancements given that approximately 98% of crashes resulted in either injuries or fatalities. The relatively small proportion of PDO incidents emphasizes the need for enhanced pedestrian safety measures. Even when crashes are non-fatal, pedestrians are frequently harmed, highlighting the critical importance of targeted interventions to reduce both injury-related and fatal outcomes.
@@ 4.2 Bus stops enviornment
Among the 596 bus stops analyzed, 280 (46.97%) locations did not report any pedestrian crashes during the study period. These locations were selected based on specific criteria outlined in the subsection 4.2. As shown in Fig. 4.2, the crash distribution is heavily skewed towards stops with zero recorded crashes. The skewness of the data, calculated at 2.87, indicates significant right skewness, primarily due to the high number of stops without any crashes.
Fig. 6.: KABCO crash distribution
@@
This study distinguishes itself by incorporating various explanatory variables, capturing spatial and temporal characteristics often excluded from the statistical model. Table 1 provides descriptive statistics for the response variables, including the crash data and other associated explanatory variables. The study draws data from multiple sources: CRIS, RHiNo, Trinity Metro, Google Earth®, and Google Street View® to ensure a comprehensive analysis. The collected variables are organized into categories covering exposure features, bus stop design characteristics, road features, and contextual elements influencing pedestrian safety. Exposure features reflect traffic and pedestrian volumes, providing insights into potential crash risk. Bus stop design characteristics describe the physical attributes of the stops, including their design and proximity to intersections. At the same time, road features capture elements of the surrounding roadway, such as lane count, speed limits, and median presence. Contextual elements account for additional factors, such as lighting and land use, that may also affect safety outcomes.
To capture the severity of crashes in a nuanced way, the crash data are classified into three groups. The first group includes total crashes (KABCO), encompassing all pedestrian-involved crashes. The second group focuses on KABC crashes, comprising incidents that resulted in fatalities (K), incapacitating injuries (A), or other injuries (B and C). The third group includes KAB crashes, concentrating on severe incidents involving fatalities and injuries classified as K, A, or B. This classification framework allows for a more detailed investigation into the relationship between crash severity and explanatory variables.
Table 1. presents the descriptive statistics for these variables, highlighting key aspects of the bus stop environments. The mean number of crashes per bus stop is 0.752, with a maximum of 8 crashes recorded at a bus stop. KABC crashes have a lower mean of 0.526, with a maximum of 6 crashes, while KAB crashes exhibit a mean of 0.314, with a maximum of 5 crashes. These numbers emphasize the variability in crash occurrences across the study area.
The study also includes several continuous variables, such as Average Annual Daily Traffic (AADT), which captures traffic exposure near bus stops. The distribution of AADT ranges from 166 to 42056 vehicles/day, with a mean of 13558.2 vehicles. While active pedestrian activity was not included, boarding and alighting volumes may offer insights into pedestrian activity at each stop. The nearest intersection distance from bus stops, ranging from 0.4 to over 2100 feet, was included as an explanatory variable to assess spatial relationships with pedestrian crashes. Other continuous variables considered in this study include median width, roadway speed limit at the bus stop location, number of lanes, presence of schools, number of parks within half a mile of the bus stops, and the number of nearby stops within a quarter mile. The dataset used in this study is supplemented with categorical variables such as the type of nearest intersections, presence of crosswalks and sidewalks, curve and median types, lighting conditions, area characteristics, bus stop shelter design, proximity of intersections to bus stops, and overall bus stop design, all of which may affect pedestrian crash risk.
@@
AADT Traffic Volume on the street near bus stop 166 42056 13558.2 8828.04
Avg on Average boarding 0 205.4 11.93 19.79
Avg off Average alighting 0 215.2 15.24 23.15
Dis to int Distance between bus stop and nearest intersection corner 0.4 2106 180.60 206.29
@@
Speed limit Posted speed limit 20 65 36.63 7.95
@@
School Number of schools within a half-mile walk of the bus stop 0 4 0.70 0.92
Park count Number of parks within a half-mile walk of the bus stop 0 6 0.55 0.81
@@
Int t Nearest intersection type Signalized (1) 384 237
Non-signalized (0) 212 206
Marked Xwalk Presence of marked crosswalk within 250ft of bus stop No (1) 223 268
Yes (0) 373 175
Med t Median types Undivided (1) 299 267
@@
Area Area type Com (0) 332 280
Res (1) 78 55
Mix (2) 186 108
@@
Curve Presence of curve near bus stop Yes (1) 55 40
No (0) 474 245
@@
Proximity The proximity of intersection from bus stop location Far (0) 306 210
Near (1) 206 166
Midblock (2) 84 70

### v4

@@ Abstract
Pedestrian safety remains a pressing concern near bus stops along urban transit, where frequent pedestrian-vehicle interactions occur. While prior research has primarily focused on intersections and midblock locations, bus stops have often been treated as secondary contributors rather than as distinct sites requiring targeted safety assessments. This has left a critical gap in understanding how traffic exposure, roadway characteristics, and bus stop design features specifically influence pedetrain crash risks around bus stop locations. To address these gaps, this study develops a comprehensive framework focused on pedestrian safety in the vicinity of bus stops. The proposed approach employs a Random Parameters Negative Binomial-Lindley (RPNB-L) model to account for unobserved heterogeneity and site-specific variability. Using data from 596 bus stops in Fort Worth, Texas (2018–2022), the model identifies that higher pedestrian crash frequencies are significantly associated with increased AADT, elevated boarding activity, and the absence of key safety elements such as crosswalks, medians, and lighting. Conversely, far-side bus stop placement, signalized intersections, sidewalks, and mixed-use development are associated with lower crash risks. Roads near schools and those with speed limits of $\leq$35 mph show elevated crash risk. To support proactive safety management, the study integrates a Full Bayes-based Potential for Safety Improvement (PSI) metric, enabling the identification of hazardous stops and high-risk corridors. By unifying advanced count-based modeling with strategic risk prioritization, this research offers actionable, data-driven insights for improving pedestrian safety near bus stops.
@@ 1 Introduction
Accordingly, the main objective of this study is to develop a comprehensive framework for assessing pedestrian safety near bus stops, one that explicitly accounts for traffic exposure, roadway characteristics, and bus stop design features. More specifically, this study (i) proposes a pedestrian crash analysis framework tailored to bus stops along urban transit corridors; (ii) applies a random parameters approach to capture unobserved heterogeneity driven by site-specific variations; and (iii) employs a Full Bayes (FB) based Potential for Safety Improvement (PSI) metric to identify and prioritize high-risk transit corridors proactively. By addressing notable gaps in the existing literature, this work offers a systematic approach to understanding and mitigating pedestrian crashes near bus stops.
To accomplish these goals, the study integrates five years (2018–2022) of pedestrian crash data from 596 bus stop locations in Fort Worth, Texas, with complementary datasets on roadway conditions, traffic exposure, and bus stop design characteristics. Traffic exposure data were obtained from local agency records, while roadway and design attributes were extracted using the Rhino database and Google Earth & Street View. To effectively capture site-specific variation and unobserved heterogeneity in crash frequencies, the study adopts the RPNB-L model as its core analytical method. Baseline models, including NB-L and RPNB-GE, are also estimated for comparison. Additionally, the Full Bayes-based PSI metric is applied to systematically identify hazardous bus stops and transit corridors, enabling a proactive, data-driven safety management approach. By bridging these critical gaps in the literature, this study contributes a novel, transferable framework for evaluating and improving pedestrian safety near bus stops in urban settings.
@@ 3.1 Negative Binomial-Lindley (NB-L)
$P(Y_{i}=y_{i}|\mu_{i},\phi,\theta)=\int\text{NB}(y_{i}|\phi,\lambda_{i}\mu_{i})\text{Lindley}(\lambda_{i}|\theta)\,d\lambda_{i}$ (1)
@@
Where $\mu_{i}$$\exp(b_{0}+\sum_{j=1}^{q}b_{j}X_{j})$ represents the expected number of crashes, and $E(\lambda_{i})$ = $\frac{\theta+2}{\theta(\theta+1)}$ is the expected value of the Lindley-distributed term. Replacing the values for $\mu_{i}$ and $E(\lambda_{i})$, the mean response function becomes:
@@ 4 Data description
This section outlines the data foundation of the study. Subsection 4.1 describes the five years of pedestrian crash data obtained from TxDOT’s Crash Records Information System (CRIS) database. Subsection 4.2 details the complementary data on bus stop design, roadway features, and traffic exposure. Together, these datasets enable a comprehensive and context-rich assessment of pedestrian safety near bus stops.
@@ 4.1 Pedestrian crash data
This study utilized bus stop data obtained from Trinity Metro, the public transit agency serving the Fort Worth metropolitan area. The data were accessed through the agency’s official website (https://www.trinitymetro.org) and included Geographic Information Systems (GIS) shapefiles containing spatial and attribute information for approximately 2500 bus stops. Each stop was identified by its corresponding main street and nearby cross-street and included details on amenities and wheelchair accessibility. In addition to the spatial data, Trinity Metro provided two supplementary spreadsheets reporting passenger boarding and alighting counts, disaggregated by stop location and service day type (weekday, Saturday, and Sunday). In the absence of direct pedestrian activity data, these counts served as a proxy for estimating relative crowding levels at each stop.
In alignment with the study’s objective, historical pedestrian crash data were systematically integrated with the stop dataset. The crash records for 2018 to 2022 were obtained from CRIS. This dataset included only reportable crashes, defined as incidents involving a fatality, injury, or property damage exceeding \$1000, and restricted to public roadways. To spatially associate pedestrian crashes with bus stop locations, a 250 ft buffer was established around each stop. This distance was informed by prior literature (e.g., Craig et al. 2019), which noted that the majority of pedestrian crashes related to transit access occur within 150 feet of a stop. Building on this, and through a trial-and-error procedure aimed at generating reliable vehicle-pedestrian crash concentration maps, a 250 ft buffer was determined to provide an optimal balance. The threshold was sufficiently wide to capture pedestrian movements and crossing behaviors associated with bus stop access, yet narrow enough to reduce the likelihood of including unrelated incidents on adjacent roadways. To maintain spatial precision, crashes located within the buffer but occurring on cross-streets were manually reviewed and excluded from the analysis. Given the challenges in collecting detailed geometric and operational data for all bus stops, a targeted sampling strategy was adopted to enable a comprehensive yet manageable evaluation. Stops were selected using two clearly defined criteria to represent both high and low-risk environments:
@@
    Bus stops with at least one recorded pedestrian crash.
@@
    Bus stops exhibiting reasonable pedestrian activity (as indicated by boarding and alighting data) but with no recorded crashes.
The first criterion captures stops with documented pedestrian crash history during the study period, enabling a reactive analysis of contributing factors. To ensure inclusivity across the full spectrum of crash severities, 316 bus stops with one or more pedestrian crashes were designated as ”case” stops. Among these, 75.3% had exactly one crash, 13.9% had two crashes, and a small number had three or more, with a maximum of eight at a single stop. This distribution, shown in Fig. 4.1(a & c), follows a long-tail pattern, most crash-involved stops experience low crash frequencies. Applying a stricter threshold (e.g., two or more crashes) would have excluded nearly 90% of these locations, greatly limiting sample diversity and statistical power. Including all stops with at least one crash thus ensures a more inclusive and representative crash risk analysis.
For the second criterion, a quantitative threshold was applied to define “reasonable pedestrian activity.” As illustrated in Fig. 4.1(b), most non-crash (control) stops had low boarding and alighting volumes: 1244 stops had 0–25 passengers/day, and 494 had 25–50 passengers/day. Including these would have introduced a major imbalance in exposure relative to the 316 case stops. Although higher-activity categories (50–75 and $\geq$100 passengers/day) contained fewer stops (n = 166 and 25, respectively), they still offered limited comparability. To mitigate this imbalance and adhere to resource constraints during data collection and fieldwork, the study limited control stops to those with $\geq$75 daily passengers. This threshold captured the upper end of the exposure distribution and provided a control sample reasonably aligned in size and exposure level with the case group. Although alternative thresholds (e.g., $\geq$50 or $\geq$100 passengers/day) were not tested in this study, future research could assess the sensitivity of results to these variations. The adopted threshold in this study reflects a practical balance between analytical rigor and feasibility, ensuring a matched, exposure-informed comparison between crash-prone and non-crash locations. By incorporating both types of stops, the analysis identifies overrepresented risk factors at hazardous sites and potential protective features at safer ones, strengthening the overall understanding of pedestrian safety dynamics near bus stops.
Fig. 2.: Distribution of ”Case” and ”Control” stops: (a) stops classification, (b) daily ridership at ”Control” stops, (c) frequency distribution among selected stops.
Fig. 4.1 presents a flowchart summarizing the data preparation process, including the integration of pedestrian crash records with bus stop and roadway information. Based on the sampling criteria described earlier, a total of 596 bus stops were selected for detailed analysis (Fig. 4.1). To support this analysis, pedestrian crash data were integrated with roadway geometric and traffic characteristics extracted from TxDOT’s 2019 Roadway Highway Inventory Network Offload (RHiNo) database, a comprehensive GIS-based dataset.
@@
The RHiNo database contains location-specific roadway information for both state and local roads, including system type, facility type, number of lanes, surface width, shoulder width, median type, and traffic volumes (current and historical). When key geometric features, such as the presence of the sidewalk, the availability of the crosswalk, or the lighting conditions, were missing, additional data were collected through Google Earth® and Google Street View®, which served as effective tools for field validation and remote assessment.
@@
The final dataset included 448 pedestrian crashes, categorized by severity. Among these, 46 (10.3%) were fatal (K-level), 394 (87.%) involved injuries (A, B, or C-level), and 8 crashes (1.8%) were classified as property damage only (PDO). The yearly distribution of crash severity is shown in Fig. 4.1, while Fig. 4.1 illustrates general crash characteristics, including speed limit, day of the week, lighting, severity distribution, and weather conditions.
These underscore the substantial risk to pedestrians near bus stops—with nearly 98% of crashes resulting in injuries or fatalities, the data highlight an urgent need for targeted safety improvements. The low proportion of PDO crashes further emphasizes the vulnerability of pedestrians in traffic environments, reinforcing the critical importance of interventions aimed at reducing both fatal and injury-related incidents.
@@ 4.2 Bus stops enviornment
Among the 596 bus stops analyzed, 280 (46.97%) had no recorded pedestrian crashes during the study period. These non-crash locations were selected based on the inclusion criteria described in Subsection 4.1. As illustrated in Fig. 4.1(c), the distribution of crash frequency is heavily right-skewed, with a skewness value of 2.87, confirming the dominance of non-crash stops in the dataset..
@@
This study distinguishes itself by integrating a diverse set of explanatory variables that capture spatial, geometric, and contextual dimensions often omitted in traditional modeling approaches. Table 1 presents descriptive statistics for the response variables and associated predictors. Data were compiled from multiple sources, including CRIS, RHiNo, Trinity Metro, Google Earth®, and Google Street View®, to ensure a comprehensive understanding of bus stop environments. The explanatory variables are organized into four overarching categories: exposure-related features, bus stop design characteristics, roadway attributes, and contextual elements. Exposure-related features reflect traffic and pedestrian activity near bus stops, incorporating metrics such as Average Annual Daily Traffic (AADT), boarding, and alighting counts. Bus stop design characteristics describe the physical layout and configuration of the stops, including design type and proximity to intersections. Roadway attributes capture infrastructure elements such as speed limits, lane counts, median type, and intersection distance. Contextual elements, such as land use type, lighting conditions, and nearby schools or parks, provide additional insight into the operating environment and pedestrian safety implications.
To evaluate the influence of crash severity, the crash data were stratified into three groups. The first group includes all pedestrian-involved crashes (KABCO). The second group (KABC) narrows the focus to fatal and injury crashes, excluding property damage only. The third group (KAB) concentrates on the most severe incidents, including fatalities and incapacitating or non-incapacitating injuries. This stratification enables a more nuanced investigation of the factors contributing to crash severity.
As shown in Table 1, the mean number of crashes per bus stop is 0.752, with a maximum of 8 crashes at a single location. The mean KABC crash frequency is 0.526 (maximum = 6), while the KAB category shows a mean of 0.314 (maximum = 5). These values demonstrate considerable variation in crash occurrences across the study area, underscoring the importance of site-specific analysis.
In addition to crash data, several continuous variables were examined. AADT values ranged from 166 to 42056 vehicles per day, with a mean of 13558.2, reflecting the variation in traffic exposure near bus stops. Although direct pedestrian counts were unavailable, boarding and alighting volumes served as reasonable proxies for pedestrian activity. The distance to the nearest intersection ranged from 0.4 to 2106 feet, enabling analysis of spatial relationships between crashes and network geometry. Other continuous variables include median width, posted speed limit, number of lanes, number of nearby schools and parks, and the count of adjacent transit stops within a 0.25-mile radius. Several categorical variables were also included to account for environmental and design-related influences. These include intersection type (signalized vs. non-signalized), presence of crosswalks and sidewalks, roadway curvature, lighting conditions, area type (commercial, residential, or mixed), bus stop design (curbside vs. other), proximity of the stop to intersections (near, far, or midblock), and shelter availability. These features may collectively influence pedestrian behavior, visibility, and driver response, thus affecting crash risk. Table 1 summarizes these variables, illustrating the diverse conditions under which pedestrian crashes occur and providing a foundation for subsequent modeling and risk analysis.
@@
AADT Traffic Volume on the street near bus stop 166 42056 13540.6 8827.85
Avg on Average boarding 0 769.0 53.20 66.78
Avg off Average alighting 0 831.0 56.63 60.64
Dis to int Distance between bus stop and nearest intersection corner 0.4 2106 198.94 241.04
@@
Speed limit Posted speed limit 20 65 36.63 7.78
@@
School Number of schools within a half-mile walk of the bus stop 0 6 0.82 1.12
Park count Number of parks within a half-mile walk of the bus stop 0 6 0.64 0.84
@@
Int t Nearest intersection type Signalized (1) 382 237
Non-signalized (0) 214 206
Marked Xwalk Presence of marked crosswalk within 250ft of bus stop No (1) 234 176
Yes (0) 362 267
Med t Median types Undivided (1) 311 267
@@
Area Area type Com (0) 307 216
Res (1) 118 116
Mix (2) 171 111
@@
Curve Presence of curve near bus stop Yes (1) 38 18
No (0) 558 425
@@
Proximity The proximity of intersection from bus stop location Far (1) 301 207
Near (0) 226 236
Midblock (2) 69 0